In a study using 16 samples and no indication of a normal distribution, the distribution that should be used to calculate confidence intervals is

a. a t distribution with 16 degrees of freedom

b. a t distribution with 17 degrees of freedom

c. a t distribution with 15 degrees of freedom

d. a normal distribution
Answer #1

solution:

n =16

Degrees of freedom = df = n - 1 =16 - 1 = 15

c. a t distribution with 15 degrees of freedom